OpenStudy (anonymous):

Is 0.99999999... a rational number?

OpenStudy (chihiroasleaf):

yes.., it is

OpenStudy (anonymous):

So can u express it in the form a/b where a and b are integers

OpenStudy (unklerhaukus):

one

OpenStudy (anonymous):

just for fun rational numbers irrational number

OpenStudy (anonymous):

But 0.9999999999999... must very close to one...... how is it one

OpenStudy (anonymous):

there must be a difference of 0.000000000...0001 between 1 and 0.999999...

OpenStudy (anonymous):

right?

hartnn (hartnn):

i don't think its rational 0.11111..... = 1/9 is rational, but if u multiply it by 9 to get 0.9999.... u get 9/9 = 1 instead and not 0.99999..... so can't be expressed as ratio of integers, so irrational.....

OpenStudy (anonymous):

Agree with @hartnn

OpenStudy (unklerhaukus):

\[x=0.999...\]\[10x=9.999...\]\[10x-x=9.999...-0.999...\]\[9x=9\]\[x=\frac 99=1\]

OpenStudy (chihiroasleaf):

isn't a number that has a sequence of digits repeats infinitely is a rational number?

OpenStudy (anonymous):

\[\frac{1}{3}=0.3r\] \[3\times \frac{1}{3}=3 \times 0.3r\] \[1=0.9r\] where 0.3r means 0.3 recurring etc.

OpenStudy (anonymous):

So 0.99999999 recurring =1 which is rational.

OpenStudy (anonymous):

0.99... = 1 Why? http://www.khanacademy.org/math/vi-hart/v/9-999----reasons-that--999------1

OpenStudy (chihiroasleaf):

agree with @UnkleRhaukus and @Traxter

OpenStudy (lgbasallote):

repeating decimals are rational

OpenStudy (anonymous):

So I guess 0.99... is actually a natural number as well?

OpenStudy (anonymous):

But i must add 0.00000000000000000000000000000000000000000000000000000000...1 to 0.9999999999999999999999999999999999999999999... to get 1 right?

OpenStudy (anonymous):

no. watch the video I linked :) A lot of people can't believe 0.99... = 1 because they can't believe that infinity is infinite, and they think that there is one last digit.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

But my net is too slow

OpenStudy (anonymous):

@sauravshakya It's good that you're questioning the reasoning behind this. Have a look at my proof above and you'll see why this is though. Well done for having an inquisitive mind and not just taking things as they are!

OpenStudy (anonymous):

Well, I'll try to explain it. If x = 0.99... then 10x = 9.99.. 10x-x = 9 x = 1 You must understand what happens here. The number got shifted. If it were a rational number, we couldn't do this, because: y = 0.9999999 10y = 9.9999990 But because the decimals repeat infinitely, we can do this, and it all works out.

OpenStudy (anonymous):

I mean, if the decimals weren't repeating infinitely (instead of "if it were a rational number"). Sorry

OpenStudy (anonymous):

@sauravshakya my answer is no

OpenStudy (anonymous):

Can u pLZ explain why no

OpenStudy (anonymous):

because it is repeating

OpenStudy (anonymous):

Then it must be rational

OpenStudy (mayankdevnani):

@pourushgupta you are wrong!!!

OpenStudy (mayankdevnani):

@sauravshakya it is rational no.

OpenStudy (mayankdevnani):

because it is terminating..

OpenStudy (mayankdevnani):

@sauravshakya where you are......lol

OpenStudy (pradipgr817):

@hartnn I think in 1/9 = 0.1111111....... the right hand side is just a approximate representation of the left. As right will be equal to left only when 0.11111.... series is infinite.

OpenStudy (anonymous):

0.999999..... and 1.0000000..... are both representations of 1

hartnn (hartnn):

yes, i considered ..... as infinite only ..... like 0.99999999...till infinity. which actually is 1.

OpenStudy (pradipgr817):

In a similar was we can say that 00.99999..... is an approximate representation of a rational number. which can be found using procedure given above.

OpenStudy (pradipgr817):

so 0.99999..... is rational I think.

hartnn (hartnn):

and a natural number!

OpenStudy (pradipgr817):

@hartnn... approximately. The rational number it represent is natural also

OpenStudy (anonymous):

I think 0.9999999999999999... can be written as 9/10+9/100+9/1000+9/10000+...

OpenStudy (anonymous):

where first term is 9/10 and common ratio is 1/10

hartnn (hartnn):

yup, that sum to infinity =1

OpenStudy (anonymous):

So, Sum=(9/10)/(1-1/10)=1

OpenStudy (pradipgr817):

yes, an elegant solution to problem